# Image-based Multimodal Models as Intruders: Transferable Multimodal Attacks on Video-based MLLMs

### v2

@@ B.3 Ablation study
Table 9: The results on the ActivityNet-200 for video understanding tasks. All scores range from 1 to 5, with lower scores indicating better attack performance. $\downarrow$ represents the performance drop compared to the clean video samples. * indicates a white-box attack.
B.4 Analysis
In this section, we will discuss the consistency of intermediate features between IMM and V-MLLM, as well as present more cases of successful adversarial attacks.
Discussion. To experimentally validate the effectiveness of Eq (6) in the main content, we analyze how the cosine similarity between adversarial and benign features in IMM/V-MLLMs evolves as the iteration number increases. The Pearson Correlation Coefficient (PCC) [35] is used to quantify the linear correlation between cosine similarity trends computed from both IMM and V-MLLMs. Fig. 11 presents the PCC analysis of these trends, using BLIP-2 and four different V-MLLMs. As shown, all PCC values exceed 0.90, indicating a strong positive linear relationship between the directional changes of intermediate features in IMM and V-MLLM. This suggests that perturbations in IMM’s image features can effectively disrupt the intermediate features of video samples in V-MLLMs. Notably, the PCC values between BLIP-2 and VideoChat, as well as Video-LLaMA, are exactly 1, which aligns with the highest ASR values observed for these models in Tab. 6. The slightly lower PCC values with Chat-UniVi and LLaVA-NeXT-Video correspond to the lower ASR values, demonstrating that a higher PCC between IMM and V-MLLMs indicates better adversarial transferability.
@@ B.4 Analysis
Appendix C Algorithm
The complete I2V-MLLM Attack process is described in Algorithm 1.
@@ Appendix C Algorithm
Output: The adversarial sample $V_{adv}$.

### v3

@@ B.4 Ablation study
Table 12: The results on the ActivityNet-200 for multi-faceted video understanding tasks. All scores range from 1 to 5, with lower scores indicating better attack performance. $\downarrow$ represents the score reduction compared to the clean video samples. * indicates a white-box attack.
B.6 Analysis
In this section, we will discuss the consistency of intermediate features between I-MLLM and V-MLLM, as well as present more cases of successful adversarial attacks.
Discussion. To experimentally validate the effectiveness of Eq (6) in the main content, we analyze how the cosine similarity between adversarial and benign features in I-MLLM/V-MLLMs evolves as the iteration number increases. The Pearson Correlation Coefficient (PCC) (Schmee_1986) is used to quantify the linear correlation between cosine similarity trends computed from both I-MLLM and V-MLLMs. Fig. 11 presents the PCC analysis of these trends, using BLIP-2 and four different V-MLLMs. As shown, all PCC values exceed 0.90, indicating a strong positive linear relationship between the directional changes of intermediate features in I-MLLM and V-MLLM. This suggests that perturbations in I-MLLM’s image features can effectively disrupt the intermediate features of video samples in V-MLLMs. Notably, the PCC values between BLIP-2 and VideoChat, as well as Video-LLaMA, are exactly 1, which aligns with the highest ASR values observed for these models in Tab. 2 in the main content. The slightly lower PCC values with Chat-UniVi and LLaVA-NeXT-Video correspond to the lower ASR values, demonstrating that a higher PCC between I-MLLM and V-MLLMs indicates better adversarial transferability.
@@ B.6 Analysis
Output: Adversarial sample $V_{adv}$.
@@
Appendix C Algorithm
The complete I2V-MLLM Attack process is described in Algorithm 1.
Figure 13: Comparison of adversarial video frames generated by I2V-MLLM under different perturbation bounds.
Figure 14: Comparison of adversarial video frames generated by I2V-MLLM under different perturbation bounds.